Could someone help me please. I’m really bad at Computer Science Engineering :(
Sheet:flights! New York Boston New York New YorkChicag American Continen at least 1 fly to least Sheet:analysis! Sheet:costs! 500 New York Within Budget Fare 2 Airline 4 American 5 Continental 600 over budget 420.00 360 within budget S 288.00 You are going to complete an analysis of several airline carriers to consider entering into negotiations for a corporate discount. On sheet flights! you have compiled information regarding the number of flights from Columbus to certain key cities to which your employees frequent travel. On sheet analysis! you will be analyzing this information. Sheet costs! contains data on roundtrip airfares from Columbus to New York. You will be using this worksheet to compare these costs with your budget and to calculate possible discounted fares. Remember when referring to cells on different worksheets (other than where you are currently inputting your formula), you will need to use the sheet name in your cell address Write a formula in cell analysis!B3, which can be copied across the row and down the column into range B3:D6, to determine (True/False) if this airline has any flights to this destination. Write a formula in cell analysis!E3, which can be copied down the column, to determine (True/False) if this airline has flights to all 3 locations. Use a Boolean function 3. Write a formula in cell analysis!F3, which can be copied down the column, to determine (True/False) if this airline has flights to any of the 3 locations. Use a Boolean function.
media%2F84a%2F84aedb0d-f331-4301-a83e-b0
Answer #1

1. =IF(AND(Flights!B3:D6<>0),"TRUE","FALSE")

2. all locations

=IF(OR(B3="TRUE",C3="TRUE",D3="TRUE"),"TRUE","FALSE")

3. Atleast one location

=IF(OR(C3="TRUE",D3="TRUE",E3="TRUE"),"TRUE","FALSE")

Does not fly to boston

=IF(Flights!B3=0,"TRUE","FALSE")

4. =IF(Flights!B3=0,"TRUE","FALSE")

5. =IF(AND(B3="TRUE",C3="TRUE"), "TRUE", IF(AND(C3="TRUE",D3="TRUE"), "TRUE", IF(AND(B3="TRUE",D3="TRUE"), "TRUE", "FALSE")))

7. =IF(AND(OR(B3="TRUE",C3="TRUE"),AND(D3="TRUE")),"TRUE","FALSE")

8. =IF(B3<=500,"Within Budget","Over Budget")

9. =B3-(B3*20)/100

10. =IF(B3>5,"Frequently",IF(B3=0,"Never","Occassionally"))

Bookl Microsoft Exce HomeInsert Page LayoutFormulas Data Review View @-厅ㄨ s Cut Σ AutoSum Times New Rom, 13 -AA -==ap -wrap Text General Copy Paste ». ▲ . % , al.o Forna tional asom ble·styles. B 1 프 田 Ξ, İ İF d Merge & Center, Insert Delete Format Sort & Find & 2Clear Filter Select Editing Clear Clipboard Font Alignment Styles Cells ЕЗ IF(B3>5,Frequently,IF(B3-0, Never,Occassionally)) # of Flights from Columbus; Boston New York Chicago Flight frequently from Columbus Boston Airline New York Chicago ty Occassionally Occassionally Occassionally 3 America West 4 American 5 Continental 6 Delta 5 OccassionallyFr 4 Neven 01Occassionallv |Occassionally! Never 0 Neve Never Never 12 15 17 19 21 44トト! 100% ;